A speculator who tries to make a short-term gain on Forex trade bought a "put option" for $ 1,500 to transact EUR 1M @ 1.25, right after 30 days. However, on the date of maturity of the contract, in the open market, each EUR is worth $1.21.calculate the net gain/loss, if any , the speculator is going to earn/sustain on this arrangement
Answer #1

We are given,

Put option premium = $1,500

Spot rate = $1.25/euro

Amunt in $ = 1M*1.25 = $1.25M

Rate after 30 days = $1.21/euro

Amount to be paid in $ = 1M * 1.21 = 1.21M

The speculator bought the put option because he thought the $ will depreciate in relation to euro. But instead it appreciated and the rate after 30 days was $1.21/euro. Hence the put option expires worthless.

Speculator's

Gain = 1.25M - 1.21M = $40,000

Cost of put option = $1,500

Net Gain = 40,000 - 1,500 = $38,500

Hence the speculator will gain a net of $38,500.

If you have any doubts please let me know in the comments. Please give a positive rating if the answer is helpful to you. Thanks.